If 10% of X is equal to 25% of Y and Y is = 16, what is the value of X?
Answer by stanbon(75887)

If 10% of X is equal to 25% of Y and Y is = 16, what is the value of X?
--------
0.10*X = 0.25*Y
------
0.1X = 0.25*16
0.1X = 4
X = 4/0.1
---
X = 40
--------------
Cheers,
Stan H.
----------------